# A GNN-based Spectral Filtering Mechanism for Imbalance Classification in Network Digital Twin

### v1

@@ 2.1 Graph Neural Network
The GNN can transform the graph structure data into standard representations, making them suitable for input into neural networks for training. We employ the concept of class spectral filtering introduced in [55] and [56] to enhance the learning process. This approach allows the GNN to efficiently propagate the information from the nodes and edges to their neighboring nodes. To achieve this, we use the neighbor aggregation approach in [57]. The network infrastructure is modeled as a graph $G=(V,E)$, where $V$ represents the set of nodes (network components) and $E$ represents the set of edges (connections). Each node $v_{i}$ in the graph is associated with a feature vector $x_{i}$, which captures relevant information about the corresponding network component. The graph can be represented using an adjacency matrix $A$, where $A_{ij}=1$ if a connection exists between nodes $v_{i}$ and $v_{j}$, and $A_{ij}=0$ otherwise. Node features are stored in the feature matrix $X$, where $X$ has the dimensions $|V|\times D$, with $|V|$ represents the number of nodes, and $D$ denotes the dimensionality of the feature vectors $X=[x_{1},x_{2},\ldots,x_{|V|}]$.
Many network components depend upon one another due to their interconnectivity [17]. The GNN architecture consists of multiple layers, each processing information from neighboring vertices to extract hierarchical representations. We let $h_{i}^{(l)}$ denote the representation of vertex $v_{i}$ at layer $l$ of the GNN. Information is aggregated from the neighboring vertices and transformed using learnable parameters. The update equation for $h_{i}^{(l)}$ can be expressed as follow:
@@
One of the biggest challenges is creating network digital twin datasets that accurately replicate the real-world failure scenario. This difficulty arises to several factors, such as the limited availability of real-world system data and the inherent complexity of simulating diverse failure types.
@@ 2.2 Graph Fourier Transform
We considered an undirected weighted graph $G=(V,E,w)$, where $V=\{0,1,\ldots,N-1\}$ is the vertex set, $E$ is the set of edges, and $w(i,j)$ is the weight function satisfying $w(i,j)=0$ for any $(i,j)$notin $E$. We assume all graphs are simple, meaning they have no loops or multiple edges.
Three matrices associated with $G$ are important: the adjacency matrix $W$, a degree matrix $D$, and a Laplacian matrix $L=D-W$. The Laplacian matrix is essential for Graph Fourier Transforms (GFTs).
@@
For $k=0,\ldots,N-1$, orthonormality implies that the sum of products of corresponding eigenfunctions is equal to the Kronecker delta function $\delta(i,j)$, where $\delta(i,j)$ equals 1 if $i=j$ and is zero otherwise. This study assumes that eigenvalues are arranged in ascending order, such that $\lambda_{0}\leq\ldots\leq\lambda_{N-1}$. In addition, $\lambda_{0}$ is strictly zero because the sum of rows in $L$ equals zero. The spectrum of the matrix, denoted as $\{\lambda_{k}\}_{k=0}^{N-1}$, is represented as $\sigma(L)$. The GFT in [58] of a graph feature $f:V\rightarrow\mathbb{R}$ is defined as $\hat{f}:\sigma(L)\rightarrow\mathbb{C}$, where $\sigma(L)$ represents the spectrum of the graph Laplacian $L$. It is expressed as follows:
@@
for $k=0,\ldots,N-1$. where, $u_{k}$ denotes the orthonormal eigenfunctions of $L$, and the GFT represents a feature expansion using these eigenfunctions. The inverse GFT is given by,
@@
which reconstruct the original function $f(i)$ from its spectral components. Interestingly, the GFT’s basic function in spectral graph research is highlighted by the fact that it is equal to the Discrete Fourier Transform (DFT) on cycle graphs.
@@ 2.3 Digital Twin Spectral Filtering
where, $\delta$ denotes the Kronecker delta function. The graphs $G_{1}$ and $G_{2}$ are referred to as the factor graphs of $G_{1}\square G_{2}$.
It has been established in [59] that the adjacency, degree, and Laplacian matrices of Cartesian product graph can be derived form those of its factor graphs. Two factor graphs, $G_{1}$ and $G_{2}$, with vertex sets $V_{1}=\{0,1,\ldots,N_{1}-1\}$ and $V_{2}=\{0,1,\ldots,N_{2}-1\}$ respectively are considered. Each with adjacency matrix $W_{1}$ and $W_{2}$, degree matrix $D_{1}$ and $D_{2}$, and Laplacian matrix $L_{1}$ and $L_{2}$. When the vertices of the Cartesian product graph are ordered lexicographically, such as $(0,0),(0,1),(0,2),\ldots,(N_{1}-1,N_{2}-1)$, the adjacency, degree, and Laplacian matrices of $G_{1}\square G_{2}$ can be expressed as $W_{1}\oplus W_{2}$, $D_{1}\oplus D_{2}$, and $L_{1}\oplus L_{2}$, respectively, where operator $\oplus$ denotes the Kronecker sum.
Definition 1: The Kronecker sum is defined by $A\oplus B=A\otimes I_{n}+I_{m}\otimes B$ for matrices $A\in\mathbb{R}^{m\times m}$ and $B\in\mathbb{R}^{n\times n}$, where $I_{n}$ represents the identity matrix of size $n$.
@@
This property holds for any $k_{1}=0,\ldots,N_{1}-1$ and $k_{2}=0,\ldots,N_{2}-1$. The resulting eigenvalues $\{\lambda^{(1)}_{k}+\lambda^{(2)}_{k}\}$ are nonnegative, and the corresponding eigenfunctions $\{u^{(1)}_{k}\otimes u^{(2)}_{k}\}$ are orthonormal. These conclusions can be straightforwardly derived from the fundamental properties of the Kronecker product.
@@
The Cartesian product graph is represented as $G_{1}\times G_{2}$, where $G_{n}$ for $n=1,2$ is an undirected weighted graph with a vertex set $V_{n}=\{0,1,\dots,N_{n}-1\}$. Assuming its graph Laplacian $L_{n}$ has eigenvalues $\{\lambda_{k}^{(n)}\}_{k=0}^{N_{n}-1}$ and corresponding orthonormal eigenfunctions $\{u_{k}^{(n)}\}_{k=0}^{N_{n}-1}$. The GFT of a graph feature $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on the product graph $G_{1}\times G_{2}$ is represented by:
@@
Thus, we interpret features on a Cartesian product graph as 2D features and propose a Twin GFT, enabling multiclass failure classification in Network Digital Twin environments.
Definition 1 The Twin Graph Fourier Transform (Twin-GFT) of a function $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on a Cartesian product graph $G_{1}\square G_{2}$ is a spectral representation $\hat{f}:\sigma(L_{1})\times\sigma(L_{2})\rightarrow\mathbb{C}$ defined as follows:
@@
The twin-GFT can be represented as a series of matrix-matrix multiplications. Using $N_{1}\times N_{2}$ matrices $F=(f(i_{1},i_{2}))_{i_{1},i_{2}}$ and $\hat{F}=(\hat{f}(\lambda_{k_{11}}^{(1)}+\lambda_{k_{22}}^{(2)}))_{k_{11},k_{22}}$, the twin-GFT applied to the features $f$ is expressed as follows:
@@
where $U_{n}$ denotes an $N_{n}\times N_{n}$ unitary matrix with $(i,k)$-th element $u_{k}^{(n)}(i)$ for $n=1,2$. Then its inverse transform is given by:
@@
The twin GFT has connections to existing transformations. First, when both factor graphs are cycle graphs, the twin GFT can be equivalent to the 2D-Discrete Fourier Transform and the 2D-GFT. Second, when a factor graph is a cycle graph, some cycles might be nested within other cycles, creating a complex network structure. The twin GFT is known as a joint graph and temporal Fourier transform which generalizes these existing transformations.
3 PROPOSED CF-GNN
Class-Fourier graph Neural Networks (CF-GNNs) perform well with both homophilic and heterophilic graphs and have shown good expressive capacities. However, current methods often lack adaptability. In particular, conventional approaches use a single, globally shared spectral filter $\hat{g}$, trained across the entire graph, with fixed frequency coefficients $\{\gamma_{k}\}_{k=0}^{K}$. Node-specific changes are not considered in the global filter, which is unable to distinguish between the distinct local structures connected to each node during the filtering process. Although polynomial-parameterized spectral filter learning offers a certain level of localization, it might not adequately capture subtle and diverse local structural patterns. It is intuitively reasonable to train a class-specific filter $\hat{g}_{i}(\lambda_{l})$ rather than depending on a globally consistent spectral filter $\hat{g}(\lambda_{l})$. A more efficient method is provided by $\hat{g}_{i}(\lambda_{l})$, tailored for each node $i$ to adaptively capture its local patterns. To mitigate this limitation, we revisit globally consistent spectral graph transforms in this section and propose a spectral filter learning strategy in the context of CF-GNNs.
@@ 3.1 Class-Fourier Transform Filtering
We present an adaptive localized spectral filtering method designed for Class-Fourier Transform Filtering on graph $G$, which the generalized translation operator inspires in Graph Signal Processing (GSP) in [60]. By leveraging polynomial-parameterized spectral filtering from [61], this technique ensures a more accurate and flexible representation of local patterns by fully accounting for the unique influence of the node at which the filter is applied. We introduce a generalized translation operator $T_{i}$ for any signal $g\in\mathbb{R}^{n}$ defined on a given graph $G$ and any $i\in\{0,1,\dots,n-1\}$: $T_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$
Using a Kronecker delta function in (17) and (18) $\delta_{i}$ centered at the $i$-th node $v_{i}$, we express generalized convolution as:
$T_{i}(g):=\sqrt{N}\,(g\ast\delta_{i})=\sqrt{N}\sum_{l=1}^{n}u_{l}u_{l}^{\top}(i)\hat{g}(\lambda_{l})$ (23)
where $u_{l}^{\top}(i)$ denotes the $i$-th element of the eigenvector $u_{l}$ corresponding to eigenvalue $\lambda_{l}$, and $\hat{g}(\lambda_{l})$ represents the spectral domain representation of the signal $g$.
In the context of Class-Fourier Transform Filtering, adaptive local filtering is achieved by first centering the filter signal $g$ at the target node $v_{i}$ using the generalized translation operator $T_{i}$, and then performing spectral convolution with $x$ as follows:
@@
where $x^{(\lambda_{l})}$ is the spectral domain representation of the input signal $x$, and $u_{l}^{\top}(i)$ is the $i$-th element of the eigenvector $u_{l}$, which corresponds to the eigenvalue $\lambda_{l}$.
Establishing that: $\hat{g}_{i}(\lambda_{l})=\sqrt{N}\,u_{l}(i)\,\hat{g}(\lambda_{l})$ in 24, the equation above becomes:
@@
As evident from the definition of $\hat{g}_{i}(\lambda_{l})$, $\hat{g}_{i}$ is a function associated with $U$. Thus, by approximating $U$, we can further establish the relationship between $\hat{g}_{i}$ and $x$. Recall the Fourier transform of an inverted graph, which allows us to deduce that $x_{i}=U_{i:}\cdot\hat{x},$ where $U_{i:}$ denotes $U$’s $i$-th row. Therefore, $U_{i:}$ can be approximately from $x_{i}$ using $U_{i:}\approx x_{i}q,$where $\hat{x}$’s pseudoinverse is $q=\text{pinv}(\hat{x})\in\mathbb{R}^{n}$. Meanwhile, take notice that $u_{l}(i)=\tilde{u}_{l}(i)$, which implies that $\hat{g}_{i}(\lambda_{l})$ can be approximately expressed as $\hat{g}_{i}(\lambda_{l})\approx\ldots$.
$\hat{g}_{i}(\lambda_{l})\approx\tilde{g}_{i}(\lambda_{l})=\sqrt{N}(x_{i}q_{l})\hat{g}(\lambda_{l})$ (26)
Where the $q_{l}$ denotes the $l$-th element of $\mathbf{q}$. Consequently, a particular filter $\tilde{g}_{i}$ corresponding to node $v_{i}$, can approximate $\hat{g}_{i}$ while simultaneously incorporating the influence of feature $x_{i}$ associated with to node $v_{i}$ in the filter coefficient estimation process.
Without loss of generality, the $K$-order polynomial approximation $\hat{P}_{K}(\cdot)$ can be utilized to directly parameterize $\tilde{g}_{i}$ without sacrificing generality; that is,$\tilde{g}_{i}=\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda),$ where $\eta_{i,k}$ represents the coefficient of $\hat{p}_{k}(\Lambda)$ for the $\tilde{g}_{i}$ filter. We only pay attention to the convolution value of the filter placed at each node is our focus, i.e.,$z_{i}=\delta_{i}(U\tilde{g}_{i}U^{\top}x).$ In this case, we utilize $\delta_{i}$ as an indicator vector $\delta_{i}=[0,\ldots,1_{i},\ldots,0]\in\mathbb{R}^{n},$ which represents a row vector with only the $i$-th element being 1 and the remaining elements being zeros. Therefore, the following is a formulation of class-oriented filtering:
$z_{i}=\delta_{i}\big{(}U\tilde{g}_{i}U^{\top}x\big{)}=\delta_{i}U\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda)U^{\top}x$
@@
Where the trainable coefficient matrix is $\Psi=[\eta_{i,k}]\in\mathbb{R}^{n\times(K+1)}$ In practice,
@@
Eq (28) can still be applied to the feature matrix $X$.

### v2

@@ 2.1 Graph neural network
The GNN can transform the graph structure data into standard representations, making them suitable for input into neural training. We employed the concept of class spectral filtering introduced in [49] and [50] to enhance the learning process. This approach allows the GNN to propagate the information efficiently from the nodes and edges to their neighboring nodes. We use the neighbor aggregation approach in [51] to achieve this propagation. The network infrastructure as depicted in Figure 3 is modeled as a graph $G=(V,E)$, where $V$ represents the node set (AMF, UDM, AUSF) and $E$ represents edge set. Each node $v_{i}$ in the graph is associated with a feature vector $x_{i}$, capturing relevant information regarding the corresponding network component. The graph can be represented using an adjacency matrix $A$, where $A_{ij}=1$ if a connection exists between nodes $v_{i}$ and $v_{j}$, and $A_{ij}=0$ otherwise. Node features are stored in the feature matrix $X$, where $X$ has the dimensions $|V|\times D$, where $|V|$ represents the number of nodes, and $D$ denotes the dimensionality of the feature vectors $X=[x_{1},x_{2},\ldots,x_{|V|}]$.
Many network components depend upon one another due to their interconnectivity [52]. The GNN architecture comprises of multiple layers, each processing information from neighboring vertices to extract hierarchical representations. We let $h_{i}^{(l)}$ denote the representation of vertex $v_{i}$ at layer $l$ of the GNN. Information is aggregated from the neighboring vertices and transformed using learnable parameters. The update equation for $h_{i}^{(l)}$ can be expressed as follows:
@@
One of the greatest challenges is creating network digital twin datasets that accurately replicate the real-world failure scenarios. This problem arises from several difficulties, including the limited availability of real-world system data and the inherent complexity of simulating various types of failure.
@@ 2.2 Graph Fourier transform
This work considers an undirected weighted graph $G=(V,E,w)$, where $V=\{0,1,\ldots,N-1\}$ denotes the vertex set, $E$ represents the edge set, and $w(i,j)$ indicate the weight function satisfying $w(i,j)=0$ for any $(i,j)$notin $E$. This graph is assumed to be simple, meaning it contains no self-loops. Each node in this graph corresponds to a functional entity in the 5G core architecture, and edges represent measurable interactions (e.g., packet transfer, registration signaling, GTP flows). These matrices associated with $G$ are important: the adjacency matrix $W$, a degree matrix $D$, and a Laplacian matrix $L=D-W$. The Laplacian matrix is essential for GFTs.
@@
For $k=0,\ldots,N-1$, orthonormality implies that the sum of the products of the corresponding eigenfunctions equal the Kronecker delta function $\delta(i,j)$, where $\delta(i,j)$ equals 1 if $i=j$ and is zero otherwise. This study assumes that eigenvalues are arranged in ascending order, such that $\lambda_{0}\leq\ldots\leq\lambda_{N-1}$. In addition, $\lambda_{0}$ is strictly zero because the sum of rows in $L$ equals zero. The matrix spectrum, denoted by $\{\lambda_{k}\}_{k=0}^{N-1}$, is represented as $\sigma(L)$. The GFT in [53] of a graph feature $f:V\rightarrow\mathbb{R}$ is defined as $\hat{f}:\sigma(L)\rightarrow\mathbb{C}$, where $\sigma(L)$ represents the spectrum of the graph Laplacian $L$. It is expressed as follows:
@@
for $k=0,\ldots,N-1$, where, $u_{k}$ denotes the orthonormal eigenfunctions of $L$, and the GFT represents a feature expansion using these eigenfunctions. The inverse GFT is given by,
@@
which reconstructs the original function $f(i)$ from its spectral components. The basic function of the GFT in spectral graph research is highlighted because it is equal to the discrete Fourier transform on cycle graphs.
@@ 2.3 Digital twin spectral filtering
where $\delta$ denotes the Kronecker delta function. The graphs $G_{1}$ and $G_{2}$ are referred to as the factor graphs of $G_{1}\square G_{2}$.
The adjacency, degree, and Laplacian matrices of the Cartesian product graph can be derived from those of its factor graphs [54]. Two factor graphs, $G_{1}$ and $G_{2}$, with vertex sets $V_{1}=\{0,1,\ldots,N_{1}-1\}$ and $V_{2}=\{0,1,\ldots,N_{2}-1\}$ respectively are considered. Each with adjacency matrix $W_{1}$ and $W_{2}$, degree matrix $D_{1}$ and $D_{2}$, and Laplacian matrix $L_{1}$ and $L_{2}$ respectively. When the vertices of the Cartesian product graph are ordered lexicographically, such as $(0,0),(0,1),(0,2),\ldots,(N_{1}-1,N_{2}-1)$, the adjacency, degree, and Laplacian matrices of $G_{1}\square G_{2}$ can be expressed as $W_{1}\oplus W_{2}$, $D_{1}\oplus D_{2}$, and $L_{1}\oplus L_{2}$, respectively, where the operator $\oplus$ denotes the Kronecker sum.
Definition 1: The Kronecker sum is defined by $A\oplus B=A\otimes I_{n}+I_{m}\otimes B$ for matrices $A\in\mathbb{R}^{m\times m}$ and $B\in\mathbb{R}^{n\times n}$, where $I_{n}$ represents an identity matrix of size $n$.
@@
This property holds for any $k_{1}=0,\ldots,N_{1}-1$ and $k_{2}=0,\ldots,N_{2}-1$. The resulting eigenvalues $\{\lambda^{(1)}_{k}+\lambda^{(2)}_{k}\}$ are nonnegative, and the corresponding eigenfunctions $\{u^{(1)}_{k}\otimes u^{(2)}_{k}\}$ are orthonormal. These conclusions can be directly derived from the fundamental properties of the Kronecker product.
@@
The Cartesian product graph is represented as $G_{1}\times G_{2}$, where $G_{n}$ for $n=1,2,G_{n}$ is an undirected weighted graph with a vertex set $V_{n}=\{0,1,\dots,N_{n}-1\}$. We assume its graph Laplacian $L_{n}$ has eigenvalues $\{\lambda_{k}^{(n)}\}_{k=0}^{N_{n}-1}$ and corresponding orthonormal eigenfunctions $\{u_{k}^{(n)}\}_{k=0}^{N_{n}-1}$. The GFT of a graph feature $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on the product graph $G_{1}\times G_{2}$ is represented by,
@@
Thus, this work interprets features on a Cartesian product graph as two-dimensional features and proposes a twin GFT, enabling multiclass failure classification in Network Digital Twin environments.
Definition 1 The twin GFT of a function $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on a Cartesian product graph $G_{1}\square G_{2}$ is a spectral representation $\hat{f}:\sigma(L_{1})\times\sigma(L_{2})\rightarrow\mathbb{C}$ defined as follows:
@@
The twin GFT can be represented as a series of matrix-to-matrix multiplications. The twin GFT applied to features f is expressed using $N_{1}\times N_{2}$ matrices $F=(f(i_{1},i_{2}))_{i_{1},i_{2}}$ and $\hat{F}=(\hat{f}(\lambda_{k_{11}}^{(1)}+\lambda_{k_{22}}^{(2)}))_{k_{11},k_{22}}$, as follows:
@@
where $U_{n}$ denotes an $N_{n}\times N_{n}$ unitary matrix with the $(i,k)$-th element $u_{k}^{(n)}(i)$ for $n=1,2$. Then, its inverse transform is given as follows:
@@
The twin GFT has connections to existing transformations. First, when both factor graphs are cycle graphs, the twin GFT can be equivalent to the 2D-Discrete Fourier Transform and the 2D-discrete Fourier transform and 2D GFT. Second, when a factor graph is a cycle graph, some cycles might be nested within other cycles, creating a complex network structure. The twin GFT is a joint graph and temporal Fourier transform, generalizing the existing transformations.
3 Proposed Class-Fourier Graph Neural Network
The CF-GNN performs well with homophilic and heterophilic graphs and has demonstrated good expressive capacities. However, the current methods often lack adaptability. Conventional approaches apply a single, globally shared spectral filter $\hat{g}$, trained across the entire graph, with fixed frequency coefficients $\{\gamma_{k}\}_{k=0}^{K}$. Node-specific changes are not considered in the global filter, which cannot distinguish between the distinct local structures connected to each node during the filtering process. Although polynomial-parameterized spectral filter learning offers a certain level of localization, it might not adequately capture subtle and diverse local structural patterns. It is intuitively reasonable to train a class-specific filter $\hat{g}_{i}(\lambda_{l})$ rather than to depend on a globally consistent spectral filter $\hat{g}(\lambda_{l})$. A more efficient method is provided by $\hat{g}_{i}(\lambda_{l})$, tailored for each node $i$ to adaptively capture its local patterns adaptively. This section revisits globally consistent spectral graph transforms and proposes a spectral filter learning strategy in the context of CF-GNNs to mitigate this limitation.
@@ 3.1 Class-Fourier transform filtering
This section introduces an adaptive localized spectral filtering method for CFT filtering on a graph $G$, inspired by the generalized translation operator from graph signal processing (GSP) in [21]. Using polynomial-parameterized spectral filtering as described in [55], this approach ensures accurate and flexible representation of local patterns by incorporating the influence of the node at which the filter is applied.
We define a generalized translation operator $T_{i}$ for any signal $g\in\mathbb{R}^{n}$ defined on a graph $G$, and for any node $i\in\{0,1,\dots,n-1\}$, as:
$T_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n},$ (23)
and specifically,
$T_{i}(g):=\sqrt{N}\,(g\ast\delta_{i})=\sqrt{N}\sum_{l=1}^{n}u_{l}u_{l}^{\top}(i)\hat{g}(\lambda_{l}),$ (24)
where $u_{l}^{\top}(i)$ denotes the $i$-th entry of the eigenvector $u_{l}$ corresponding to eigenvalue $\lambda_{l}$, and $\hat{g}(\lambda_{l})$ is the spectral representation of signal $g$.
In the context of CFT filtering, adaptive local filtering is achieved by centering the filter signal $g$ at the target node $v_{i}$ via $T_{i}$, followed by spectral convolution with $x$:
@@
Since $\hat{g}_{i}(\lambda_{l})=\sqrt{N}\,u_{l}(i)\,\hat{g}(\lambda_{l})$, we can rewrite Equation 25 as:
@@
Noting that $x_{i}=U_{i:}\cdot\hat{x}$, where $U_{i:}$ is the $i$-th row of the eigenvector matrix $U$, and $q=\text{pinv}(\hat{x})\in\mathbb{R}^{n}$, we approximate:
$U_{i:}\approx x_{i}q,\quad\hat{g}_{i}(\lambda_{l})\approx\tilde{g}_{i}(\lambda_{l})=\sqrt{N}(x_{i}q_{l})\hat{g}(\lambda_{l})$ (27)
with $q_{l}$ the $l$-th element of $q$.
This leads to a node-specific filter $\tilde{g}_{i}$ that reflects the influence of feature $x_{i}$. The filter can be parameterized using a $K$-order polynomial approximation:
$\tilde{g}_{i}=\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda),$ (28)
where $\eta_{i,k}$ are trainable coefficients. Focusing on the filter output at each node, we compute:
$z_{i}=\delta_{i}(U\tilde{g}_{i}U^{\top}x)=\delta_{i}U\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda)U^{\top}x$ (29)
@@
with $\Psi=[\eta_{i,k}]\in\mathbb{R}^{n\times(K+1)}$. This extends to a feature matrix $X$ as:
@@
CF-GNN-V: Eigenvalue-based Transformation
This variant applies a transformation to the learned features in the spectral domain that emphasizes eigenvalue-related frequency components relevant to the minority class. It introduces a class-sensitive weighting function $\gamma(\lambda)$ such that:
$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\gamma(\Lambda)\hat{p}_{k}(L)X$ (32)
Here, $\gamma(\Lambda)$ is a diagonal matrix scaling frequency components to highlight class-specific spectral patterns.
Proof of CF-GNNv
The base form of graph convolution in the spectral domain is:
$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\hat{p}_{k}(L)X$ (33)
This corresponds to filtering the signal $X$ using polynomial spectral filters $\hat{p}_{k}(\Lambda)$, followed by node-wise localization using $\delta_{i}$.
In CF-GNNv, we introduce a class-sensitive transformation $\gamma(\Lambda)$ into the filtering process:
$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\gamma(\Lambda)\hat{p}_{k}(L)X$ (34)
In spectral graph theory, filtering a signal $x\in\mathbb{R}^{n}$ is represented as:
$x^{\prime}=Ug(\Lambda)U^{\top}x$ (35)
where $g(\Lambda)$ is a diagonal matrix that applies the filter in the spectral domain.
To emphasize frequency components relevant to the minority class, we define $\gamma(\lambda)$ to selectively modulate these components. Since $\gamma(\Lambda)$ and $\hat{p}_{k}(\Lambda)$ are diagonal, we can define a composite filter:
$\tilde{g}_{i,k}(\Lambda)=\gamma(\Lambda)\hat{p}_{k}(\Lambda)$ (36)
which leads to:
$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}U\tilde{g}_{i,k}(\Lambda)U^{\top}X$
$=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}U\gamma(\Lambda)\hat{p}_{k}(\Lambda)U^{\top}X$
$=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\gamma(\Lambda)\hat{p}_{k}(L)X$ (37)
This final expression shows that CF-GNN-V modifies the spectral filter to reflect class-sensitive frequency emphasis.
3.1.1 Eigenvector-aware attention
We begin with the classical spectral graph convolution expression, as shown earlier in Eq. (35):
$x^{\prime}=Ug(\Lambda)U^{\top}x$ (38)
Now, instead of only manipulating the eigenvalues $\Lambda$ using $g(\Lambda)$, we propose to modify the spectral projection basis $U$ using eigenvector-aware attention $\alpha(U)$, where $\alpha(U)\in\mathbb{R}^{n\times n}$ could be either a full matrix or a diagonal one (simpler to implement and more stable in training).
The modified filtering of the signal is:
$x^{\prime}=U\alpha(U)g(\Lambda)U^{\top}x$ (39)
This formulation applies an attention mechanism in the spectral domain on the eigenvectors, rather than just adjusting the spectral coefficients (eigenvalues). The idea is that certain eigenvectors capture structural features (e.g., localized clusters, boundary conditions) that may be more informative for classifying failure modes, especially in minority classes.
We define the composite filter with eigenvector attention as:
$\tilde{g}_{i,k}(U,\Lambda)=\alpha(U)\hat{p}_{k}(\Lambda)$ (40)
Substituting into the spectral convolution yields:
$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}U\alpha(U)\hat{p}_{k}(\Lambda)U^{\top}X=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\alpha(U)\hat{p}_{k}(L)X$ (41)
This final expression shows that CF-GNNE introduces class-aware attention over eigenvectors in the spectral filtering process, complementing the eigenvalue-focused approach of CF-GNNv.